One year ago you signed a​ 4-year lease on a car. The sticker price on the car was ​$22,000, you made no down​ payment, the lease rate was 3​%, and the monthly payments (starting at the time of​ signing) were $260.51. The buyout was $11,500 due at the end of the lease term. Now​ (one year after signing the lease but just before the 13th lease​ payment) a new car has been released with better styling and​ 20% more horsepower. You want to get out of your lease on the old car and lease one of the new ones. How much do you still​ owe?
Answer #1

The amount owed wlll be equal to the present value (at current time) of the remaining lease payment and the promised buyout amount, discounted at the lease rate of 3 %.

Lease Payments = $ 260.5, Lease Rate = 3 %, Number of Lease Payments Remaining = 36 and Buyout Payment = $ 11500

Monthly Lease Rate = (3/12) = 0.25%

Therefore, amount still owed = 260.5 x (1/0.0025) x [1-{1/(1.0025)^(36)}] x (1.0025) + 11500 / (1.0025)^(36) = $ 19491.46